OpenStudy (anonymous):

PLEASE HELP!!!!!!!!!!!! ALGEBRA 1 5-(x+5) > -2 (x+4)

OpenStudy (anonymous):

-5x+ 25 > -2x -8

OpenStudy (anonymous):

combine like terms -5+ 2= -3x+25=8 -3x =33/3 = 11

OpenStudy (anonymous):

do u want me to draw it

OpenStudy (anonymous):



OpenStudy (anonymous):

-3x>-33 x>11

OpenStudy (anonymous):

ya i meant to put that sign wasnt lookin

OpenStudy (anonymous):

good eye @AJ01

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

np

OpenStudy (anonymous):

Thanks so much you guys!!!!

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

@dancingdiamond if u need anymore help with alg 1 tag me just when u have graphs i suck at those still

OpenStudy (anonymous):

when u have graphs dont tag me

OpenStudy (anonymous):

Thanks so much!!